# MCbiF: Measuring Topological Autocorrelation in Multiscale Clusterings via 2-Parameter Persistent Homology

### v1

@@ 4 The Multiscale Clustering Bifiltration (MCbiF)
The central object of our paper is a novel bifiltration of abstract simplicial complexes that encodes cluster intersection patterns in the sequence of partitions $\theta$ across the scale $t$.
@@
Let $\theta:[t_{1},\infty)\rightarrow\Pi_{X}$ be a sequence of partitions. We define $\mathcal{M}$, the Multiscale Clustering Bifiltration (MCbiF) as a bifiltration of abstract simplicial complexes:
$\mathcal{M}:=(K^{s,t})_{t_{1}\leq s\leq t}\quad\text{where}\quad K^{s,t}:=\bigcup_{t_{1}\leq s\leq r\leq t}\,\,\bigcup_{C\in\theta(r)}\Delta C,\quad t_{1}\leq s\leq t.$ (5)
In this construction, each cluster $C$ corresponds to a (|C|1)(|C|-1)-dimensional solid simplex $\Delta C:=2^{C}$, which, by definition, contains all its lower dimensional simplices [Schindler and Barahona, 2025]. This echoes natural concepts of data clustering as information compression or lumping [Rosvall and Bergstrom, 2008, 2011, Lambiotte et al., 2014], and of clusters as equivalence classes [Brualdi, 2010]. The MCbiF then aggregates all clusters (simplices) from partition $\theta(s)$ to $\theta(t)$ through the union operators, such that a $k$-simplex $\sigma=[x_{1},\dots,x_{k+1}]\in K^{s,t}$ consists of elements that are assigned to the same cluster (at least once) in the interval $[s,t]$, i.e., $x_{1},\dots,x_{k+1}\in C$ for some cluster $C\in\theta(r)$ and $r\in[s,t]$. The bifiltration thus depends not only on the lag |ts||t-s| but also on the starting scale $s$, and captures the topological autocorrelation in the sequence of partitions, see Fig. 1.
We first show that the MCbiF is indeed a well-defined bifiltration.
@@
The proof is straightforward and can be found in Appendix A.2.
The MCbiF leads to a triangular commutative diagram where the arrows indicate inclusion maps between abstract simplicial complexes (see Fig. 1). The sequence of partitions $\theta(t)$ is encoded by the complexes $K^{t,t}$ on the diagonal of the diagram. Moving along horizontal arrows corresponds to fixing a starting scale $s$ and going forward in the sequence $\theta$, thus capturing the coarse-graining of partitions. Moving along vertical arrows corresponds to fixing an end scale $t$ and aggregating $\theta$ going backwards, capturing the fine-graining of partitions.
Remark 5.
By fixing $s:=t_{1}$ (i.e., the top row in the commutative MCbiF diagram), we recover the 1-parameter Multiscale Clustering Filtration (MCF) defined by Schindler and Barahona [2025]. MCF was designed to quantify non-hierarchies in coarse-graining sequences of partitions and thus cannot capture fine-graining. For example, a large cluster $C\in\theta(s^{\prime})$ prevents MCF from detecting cluster assignment conflicts between elements $x,y\in C$ for $t\geq s^{\prime}$, see Section 4.1. In contrast, MCbiF encodes the full topological autocorrelation contained in $\theta$ by varying both the starting scale $s$ and the lag tst-s.
Applying MPH to the bifiltration $\mathcal{M}$ at dimensions $k\leq\dim K$, for $K=K^{t_{M},t_{M}}$, leads to a triangular diagram of simplicial complexes $H_{k}(K^{s,t})$ called persistence module (see Section 3.2). This persistence module has strong algebraic properties, as stated in the following proposition.
@@
See Botnan and Lesnick [2023] for definitions and Appendix A.2 for details and a full proof.
The properties in Proposition 6 are important because they guarantee algebraic stability of the MCbiF [Bjerkevik, 2021]. In particular, the finite presentation property implies stability of the MCbiF Hilbert functions $\mathrm{HF}_{k}(s,t)$ (Eq. 4) with respect to small changes in the module [Oudot and Scoccola, 2024, Corollary 8.2.]. This justifies the use of $\mathrm{HF}_{k}(s,t)$ as simple interpretable invariants for the topological autocorrelation captured by MCbiF, as exploited in Section 4.1.
4.1 Measuring Topological Autocorrelation with MCbiF
We now show how topological autocorrelation as measured by $\mathrm{HF}_{k}(s,t)$ can be used to detect cluster-assignment conflicts. We focus on dimensions $k=0,1$, for which MPH is implemented in RIVET [Lesnick and Wright, 2015]. Note that $\mathrm{HF}_{0}(s,t)$ counts the number of connected components and $\mathrm{HF}_{1}(s,t)$ the number of 1-dimensional holes in $K^{s,t}$, see Section 3.2 for details. We show below that the computation of these invariants reveals different aspects of the non-hierarchy in the sequence of partitions.
4.1.1 Low-order Non-Hierarchy in Sequences of Partitions
Hierarchy in a sequence of partitions can be understood as a refinement of partitions captured by the partition lattice.
Definition 7 (Hierarchy).
The sequence of partitions $\theta$ is hierarchical in $[s,t]$ if we have a strict sequence of refinements: either $\theta(r_{1})\leq\theta(r_{2}),\forall r_{1},r_{2}\in[s,t]$ with $r_{1}\leq r_{2}$ (agglomerative) or $\theta(r_{1})\geq\theta(r_{2}),\,\forall r_{1},r_{2}\in[s,t]$ with $r_{1}\leq r_{2}$ (divisive). We say that $\theta$ is strictly hierarchical if it is hierarchical in $t_{1},\infty)$.
One important aspect of hierarchy is the nestedness of the clusters in the sequence.
Definition 8 (Nestedness).
We say that $\theta$ is nested in $[s,t]$ when $\forall r_{1},r_{2}\in[s,t]$, we have that $\forall C\in\theta(r_{1}),C^{\prime}\in\theta(r_{2})$, one of the sets $\setminus C^{\prime}, CC$$^{\prime}\setminus C$ or $C\cap C^{\prime}$ is empty. See [Korte and Vygen, 2012, Definition 2.12]. We say that $\theta$ is strictly nested when $\theta$ is nested in $t_{1},\infty)$.
Remark 9.
It follows directly from the definitions that a hierarchical sequence $\theta$ is always nested. However, nestedness does not necessarily imply hierarchy, as illustrated by the example in Fig 3b.
We can quantify the low-order non-hierarchy in a sequence $\theta$ through the computation of the invariant $\mathrm{HF}_{0}(s,t)$ and the associated notion of 0-conflicts defined next.
Remark 10.
Each partition $\theta(t)$ can be interpreted as an equivalence relation $\sim_{t}$ given by the property of belonging to the same cluster, i.e., $x\sim_{t}y$ if $\exists\;C\in\theta(t)$ such that $x,y\in C$ [Brualdi, 2010].
Definition 11 (0-conflict and triangle 0-conflict).
  1. a)
    We say that $\theta$ has a 0-conflict in $[s,t]$ if the subposet $\theta([s,t])$ has no maximum, i.e., $\nexists r\in[s,t]$ such that $\theta(r^{\prime})\leq\theta(r)$, $\forall r^{\prime}\in[s,t]$.
  2. b)
    We say that $\theta$ has a triangle 0-conflict in $[s,t]$ if $\exists\;x,y,z\in X$ such that $\exists r_{1},r_{2}\in[s,t]$: $x\sim_{r_{1}}y\sim_{r_{2}}z$ and $\nexists r\in[s,t]$: $x\nsim_{r}y\nsim_{r}z$.
Next, we show that all triangle 0-conflicts are also 0-conflicts. Moreover, all 0-conflicts break hierarchy and triangle 0-conflicts additionally break nestedness.
@@ 4.1.1 Low-order Non-Hierarchy in Sequences of Partitions
  1. (i)
    Every triangle 0-conflict is a 0-conflict, but the opposite is not true.
  2. (ii)
    If $\theta$ has a 0-conflict in $[s,t]$, then $\theta$ is non-hierarchical in $[s,t]$.
  3. (iii)
    If $\theta$ is either coarse- or fine-graining but non-hierarchical in $[s,t]$, then $\theta$ has a 0-conflict in $[s,t]$.
  4. (iv)
    If $\theta$ has a triangle 0-conflict in $[s,t]$, then $\theta$ is non-nested in $[s,t]$.
See Appendix A.2.1 for the simple proof. Fig. 3b illustrates a 0-conflict that is not a triangle 0-conflict, and Fig. 3c shows a triangle 0-conflict.
Remark 13.
The following proposition develops a sharp upper bound for $\mathrm{HF}_{0}$ that can be used to capture 0-conflicts.
Proposition 14.
  1. (i)
    $\mathrm{HF}_{0}(s,t)\leq\min_{r\in[s,t]}|\theta(r)|$, $\forall[s,t]\subseteq[t_{1},\infty)$.
  2. (ii)
    $\mathrm{HF}_{0}(s,t)<\min_{r\in[s,t]}|\theta(r)|$ iff $\theta$ has a 0-conflict in $[s,t]$.
  3. (iii)
    $\mathrm{HF}_{0}(s,t)=|\theta(r)|$ for $r\in[s,t]$ iff $\theta(r)$ is the maximum of the subposet $\theta([s,t])$.
See Appendix A.2.1 for a full proof. Proposition 14 shows that $\mathrm{HF}_{0}$ measures low-order non-hierarchy in $\theta$ by capturing 0-conflicts. To quantify this, we introduce a global normalised measure for the sequence $\theta$, defined as follows.
@@
  1. (i)
    If $\theta$ is hierarchical in $[s,t]$, then $\mathrm{HF}_{0}(s,t)=\min(|\theta(s)|,|\theta(t)|)$. As a special case, this implies $\mathrm{HF}_{0}(t,t)=|\theta(t)|,\forall t\geq t_{1}$.
  2. (ii)
    $\bar{c}_{0}(\theta)>0$ iff $\theta$ has a 0-conflict.
  3. (iii)
    Let $\theta$ be either coarse- or fine-graining. Then, $\bar{c}_{0}(\theta)=0$ iff $\theta$ is strictly hierarchical.
A proof can be found in Appendix A.2.1.
We can further detect triangle 0-conflicts by analysing the graph-theoretic properties of the MCbiF 1-skeleton $K_{1}^{s,t}$. Recall that the clustering coefficient $\mathrm{C}$ of a graph is defined as the ratio of the number of triangles to the number of paths of length 2 in the graph  [Luce and Perry, 1949, Newman, 2018].
Proposition 17.
See Appendix A.2.1 for a proof. Let us consider the graph generated as the disjoint union of all clusters from partitions in [s,t] as cliques. This graph is equivalent to the MCbiF 1-skeleton $K_{1}^{s,t}$. Proposition 17 shows that the clustering coefficient of this graph can be used to detect triangle 0-conflicts that are not cycles. To be able to detect triangle 0-conflicts that correspond to non-bounding cycles, we turn to the 1-dimensional homology in the next section.
4.1.2 Higher-order Inconsistencies between Clusters in Sequences of Partitions
Measuring 0-conflicts in $\theta$ is only one way of capturing non-hierarchy. An additional phenomenon that can arise in non-hierarchical sequences is higher-order inconsistencies of cluster assignments across scales. These are captured by the 1-dimensional homology groups [Schindler and Barahona, 2025] and the associated notion of 1-conflict, which we define next.
Recall the definition of $1$-cycles $Z_{1}(K^{s,t})$ and non-bounding cycles $H_{1}(K^{s,t})$ (Eq.14) summarised in Appendix C.
@@ 4.1.2 Higher-order Inconsistencies between Clusters in Sequences of Partitions
The number of distinct 1-conflicts for the interval $[s,t]$ (up to equivalence of the homology classes) is given by $\mathrm{HF}_{1}(s,t)$. We first show that 1-conflicts also lead to triangle 0-conflicts and thus break hierarchy and nestedness of $\theta$.
@@
  1. (i)
    $\mathrm{HF}_{1}(s,t)\geq 1$ iff $\theta$ has a 1-conflict in $[s,t]$.
  2. (ii)
    If $\theta$ has a 1-conflict in $[s,t]$, then it also has a triangle 0-conflict.
  3. (iii)
    If $\theta$ is hierarchical in $[s,t]$, then $\mathrm{HF}_{1}(s,t)=0$.
See Appendix A.2.1 for a proof. Proposition 19 shows that a 1-conflict is a special kind of triangle 0-conflict arising from higher-order cluster inconsistencies across scales. This is illustrated in Fig. 12 and more details are provided in Examples 42 and 43 in Appendix B, which present sequences of partitions where different 1-conflicts emerge across scales. Moreover, Proposition 19 (iii) shows that the MCbiF has a trivial 1-dimensional MPH if $\theta$ is strictly hierarchical.
@@
Remark 21.
The presence of a 1-conflict for the interval $[s,t]$ signals the fact that assigning all the elements involved in the conflict to a shared cluster would increase the consistency of the sequence. Hence, when a 1-conflict gets resolved, e.g., the corresponding homology generator dies in the MPH at $(s,t^{\prime})$, $t<t^{\prime}$, then we say that $\theta(t^{\prime})$ is a conflict-resolving partition, see Schindler and Barahona [2025]. As two illustrations, consider Example 42, where a 1-conflict emerges through the interplay of partitions at scales $t=1,2,3$ and gets resolved by $\theta(4)=\hat{1}$, and Example 43, where three different 1-conflicts first emerge between scales $t=1$ and $t=6$ and then get resolved one by one by partitions $\theta(7)$, $\theta(8)$ and $\theta(9)$.
To quantify the presence of 1-conflicts in $\theta$ we introduce an unnormalised global measure for the sequence $\theta$, as follows.
@@
Let $T$ be defined as in Definition 15. The average 1-conflict is defined as:
@@
$c_{1}(\theta)>0$ iff $\theta$ has a 1-conflict. In particular, if $\theta$ is strictly nested, then $c_{1}(\theta)=0$.
Remark 24.
Figure 2: Summary of key theoretical results and their relationships indicated by arrows. Double-headed arrows represent equivalences (iff), whereas single-headed arrows represent implications (if).

### v2

@@ 3 The Multiscale Clustering Bifiltration (MCbiF)
The central object of our paper is a novel bifiltration of abstract simplicial complexes that encodes cluster intersection patterns in the sequence of partitions $\theta$ across the scale $t$. For background on simplicial complexes and bifiltrations, see Appendix E.2.
@@
Given a sequence $\theta:[t_{1},\infty)\rightarrow\Pi_{X}$, we define the Multiscale Clustering Bifiltration (MCbiF), $\mathcal{M}$, as a bifiltration of abstract simplicial complexes:
$\mathcal{M}:=(K^{s,t})_{t_{1}\leq s\leq t}\quad\text{where}\quad K^{s,t}:=\bigcup_{t_{1}\leq s\leq r\leq t}\,\,\bigcup_{C\in\theta(r)}\Delta C.$ (1)
In this construction, each cluster $C$ corresponds to a (|C|1)(|C|-1)-dimensional solid simplex $\Delta C:=2^{C}$, which, by definition, contains all its lower-dimensional simplices (Schindler & Barahona, 2025). This echoes natural concepts of data clustering viewed as information compression or lumping (Rosvall & Bergstrom, 2008; 2011; Lambiotte et al., 2014), and of clusters as equivalence classes (Brualdi, 2010). The MCbiF then aggregates all clusters (simplices) from partition $\theta(s)$ to $\theta(t)$ through the union operators, such that a $k$-simplex $\sigma=[x_{1},\dots,x_{k+1}]\in K^{s,t}$ consists of elements that are assigned to the same cluster (at least once) in the interval $[s,t]$, i.e., $x_{1},\dots,x_{k+1}\in C$ for some cluster $C\in\theta(r),\,r\in[s,t]$. The bifiltration depends not only on the lag |ts||t-s| but also on the starting scale $s$, and captures the topological autocorrelation in the sequence of partitions, see Fig. 1. We first show that the MCbiF is indeed a well-defined bifiltration.
@@
The proof is straightforward, see Appendix B.2. The MCbiF leads to a triangular commutative diagram where arrows indicate inclusion maps between abstract simplicial complexes (Fig. 1). The sequence of partitions $\theta(t)$ is encoded by the complexes $K^{t,t}$ on the diagonal of the diagram, hence the MCbiF is a complete invariant of $\theta$. Moving along horizontal arrows corresponds to fixing a starting scale $s$ and going forward in the sequence $\theta$, thus capturing coarse-graining. Moving along vertical arrows corresponds to fixing an end scale $t$ and going backwards in $\theta$, capturing fine-graining. By fixing $s:=t_{1}$ (top row in the MCbiF diagram), we recover the 1-parameter Multiscale Clustering Filtration (MCF) defined by Schindler & Barahona (2025), see Remark 24.
Applying MPH to the bifiltration $\mathcal{M}$ at dimensions $k\leq\dim K$, for $K=K^{t_{M},t_{M}}$, leads to a triangular diagram of simplicial complexes $H_{k}(K^{s,t})$ called persistence module (see Appendix E.2). We show in Proposition 27 in Appendix B.2 that the MCbiF persistence module is pointwise finite-dimensional, finitely presentable and block-decomposable (see Botnan & Lesnick (2023) for definitions). These strong algebraic properties are important because they guarantee algebraic stability of the MCbiF (Bjerkevik, 2021). In particular, the finite presentation property implies stability of the MCbiF Hilbert functions $\mathrm{HF}_{k}(s,t)$ (see Eq. 12 in Appendix E.2) with respect to small changes in the module (Oudot & Scoccola, 2024, Corollary 8.2.). This justifies the use of $\mathrm{HF}_{k}(s,t)$ as simple interpretable invariants for the topological autocorrelation captured by in the following.
3.1 Measuring Topological Autocorrelation with MCbiF
We now show how the topological autocorrelation measured by the $\mathrm{HF}_{k}(s,t)$ of MCbiF can be used to detect cluster-assignment conflicts. We focus on dimensions $k=0,1$ in this paper, where $\mathrm{HF}_{0}(s,t)$ counts the number of connected components of $K^{s,t}$ and $\mathrm{HF}_{1}(s,t)$ is the number of 1-dimensional holes in $K^{s,t}$ (see Appendix E.2). We show below that the computation of these invariants reveals different aspects of the non-hierarchy in the sequence of partitions.
Low-order Non-Hierarchy in Sequences of Partitions
An important aspect of hierarchy is the nestedness of the clusters in the sequence. We say that $\theta$ is nested in $[s,t]$ when $\forall r_{1},r_{2}\in[s,t]$, we have that $\forall\;C\in\theta(r_{1}),C^{\prime}\in\theta(r_{2})$, one of the sets $\setminus C^{\prime}, CC$$^{\prime}\setminus C$ or $C\cap C^{\prime}$ is empty. See Korte & Vygen (2012, Definition 2.12). We say that $\theta$ is strictly nested when $\theta$ is nested in $t_{1},\infty)$. It follows directly that a hierarchical sequence $\theta$ is always nested. However, the converse is not necessarily true, as illustrated by the nested, non-hierarchical sequence in Fig. 2b. To quantify this low-order non-hierarchy in sequence $\theta$, we can compute the invariant $\mathrm{HF}_{0}(s,t)$ with its associated notion of 0-conflicts defined next. Recall that each partition $\theta(t)$ can be interpreted as an equivalence relation $\sim_{t}$ where $x\sim_{t}y$ if $\exists\;C\in\theta(t)$ such that $x,y\in C$ (Brualdi, 2010).
Definition 5 (0-conflict and triangle 0-conflict).
a) We say that $\theta$ has a 0-conflict in $[s,t]$ if the subposet $\theta([s,t])$ has no maximum, i.e., $\nexists r\in[s,t]$ such that $\theta(r^{\prime})\leq\theta(r)$, $\forall r^{\prime}\in[s,t]$. b) We say that $\theta$ has a triangle 0-conflict in $[s,t]$ if $\exists\;x,y,z\in X$ such that $\exists r_{1},r_{2}\in[s,t]$: $x\sim_{r_{1}}y\sim_{r_{2}}z$ and $\nexists r\in[s,t]$: $x\nsim_{r}y\nsim_{r}z$.
Next, we show that all triangle 0-conflicts are 0-conflicts. Moreover, all 0-conflicts break hierarchy, and triangle 0-conflicts additionally break nestedness.
@@ 3.1 Measuring Topological Autocorrelation with MCbiF
(i) Every triangle 0-conflict is a 0-conflict, but the opposite is not true. (ii) If $\theta$ has a 0-conflict in $[s,t]$, then $\theta$ is non-hierarchical in $[s,t]$. (iii) If $\theta$ is either coarse- or fine-graining but non-hierarchical in $[s,t]$, then $\theta$ has a 0-conflict in $[s,t]$. (iv) If $\theta$ has a triangle 0-conflict in $[s,t]$, then $\theta$ is non-nested in $[s,t]$.
See Appendix B.2.1 for the simple proof. Fig. 2b illustrates a 0-conflict that is not a triangle 0-conflict, and Fig. 2c shows a triangle 0-conflict. The following proposition develops a sharp upper bound for $\mathrm{HF}_{0}$ that can be used to capture 0-conflicts.
@@
(i) $\mathrm{HF}_{0}(s,t)\leq\min_{r\in[s,t]}|\theta(r)|$, $\forall[s,t]\subseteq[t_{1},\infty)$. (ii) $\mathrm{HF}_{0}(s,t)<\min_{r\in[s,t]}|\theta(r)|$ iff $\theta$ has a 0-conflict in $[s,t]$. (iii) $\mathrm{HF}_{0}(s,t)=|\theta(r)|$ for $r\in[s,t]$ iff $\theta(r)$ is the maximum of the subposet $\theta([s,t])$.
See Appendix B.2.1 for the proof. Proposition 7 shows that $\mathrm{HF}_{0}$ measures low-order non-hierarchy in $\theta$ by capturing 0-conflicts. To quantify this, we introduce a global normalised measure for $\theta$.
@@
(i) If $\theta$ is hierarchical in $[s,t]$, then $\mathrm{HF}_{0}(s,t)=\min(|\theta(s)|,|\theta(t)|)$. As a special case, this implies $\mathrm{HF}_{0}(t,t)=|\theta(t)|,\forall t\geq t_{1}$. (ii) $\bar{c}_{0}(\theta)>0$ iff $\theta$ has a 0-conflict. (iii) Let $\theta$ be either coarse- or fine-graining. Then, $\bar{c}_{0}(\theta)=0$ iff $\theta$ is strictly hierarchical.
See Appendix B.2.1 for the proof. Furthermore, we can detect triangle 0-conflicts by analysing the graph-theoretic properties of the MCbiF 1-skeleton $K_{1}^{s,t}$. Proposition 30 in Appendix B.2.1 shows that a clustering coefficient $C(K_{1}^{s,t})<1$ indicates the presence of a triangle 0-conflict.
Higher-order Inconsistencies between Clusters in Sequences of Partitions
Measuring 0-conflicts in $\theta$ is only one way of capturing non-hierarchy. An additional phenomenon that can arise in non-hierarchical sequences is higher-order inconsistencies of cluster assignments across scales. These are already captured by the 1-dimensional homology groups (Schindler & Barahona, 2025) and the associated notion of 1-conflict, which we define next. Recall the definition of $1$-cycles $Z_{1}(K^{s,t})$ and non-bounding cycles $H_{1}(K^{s,t})$, see Eq.14 in Appendix E.3.
@@
The number of distinct 1-conflicts in $[s,t]$ (up to equivalence of the homology classes) is given by $\mathrm{HF}_{1}(s,t)$. We first show that 1-conflicts also lead to triangle 0-conflicts and thus break hierarchy and nestedness of $\theta$.
@@
(i) $\mathrm{HF}_{1}(s,t)\geq 1$ iff $\theta$ has a 1-conflict in $[s,t]$. (ii) If $\theta$ has a 1-conflict in $[s,t]$, then it also has a triangle 0-conflict. (iii) If $\theta$ is hierarchical in $[s,t]$, then $\mathrm{HF}_{1}(s,t)=0$.
See Appendix B.2.1 for a proof. Proposition 11 shows that a 1-conflict is a special kind of triangle 0-conflict arising from higher-order cluster inconsistencies across scales. This is illustrated in Fig. 1 and Examples 3738 in Appendix C, which present sequences of partitions where different 1-conflicts emerge across scales. Note that Proposition 11 (iii) shows that the MCbiF has a trivial 1-dimensional MPH if $\theta$ is strictly hierarchical.
@@
The presence of a 1-conflict in $[s,t]$ signals the fact that assigning all the elements involved in the conflict to a shared cluster would increase the consistency of the sequence $\theta$. Hence, when a 1-conflict gets resolved (e.g., the corresponding homology generator dies in the MPH at $(s,t^{\prime})$, $t<t^{\prime}$), then $\theta(t^{\prime})$ is called a conflict-resolving partition  (Schindler & Barahona, 2025). This is illustrated in Example 37, where a 1-conflict gets resolved by $\theta(4)=\hat{1}$, and Example 38, where three different 1-conflicts get resolved one by one by partitions $\theta(7)$, $\theta(8)$ and $\theta(9)$.
To quantify the presence of 1-conflicts in $\theta$, we introduce an unnormalised global measure.
@@
Let $T$ be as in Definition 8. The average 1-conflict is defined as:
@@
$\bar{c}_{1}(\theta)>0$ iff $\theta$ has a 1-conflict. In particular, if $\theta$ is strictly nested, then $\bar{c}_{1}(\theta)=0$.
Figure 4 in Appendix A summarises all of our theoretical results and their relationships.